OpenStudy (anonymous):

Eddie is fishing. He has 1 1/4 ounces of weight on his line, but his bait isn’t getting to the bottom of the lake. Eddie adds another 1/2-ounce weight to the line, but it's still not enough. Finally, Eddie adds a 3/4-ounce weight and the bait sinks. How much weight does Eddie now have on his line? please help

OpenStudy (texaschic101):

1 1/4 + 1/2 + 3/4 5/4 + 1/2 + 3/4 5/4 + 2/4 + 3/4 = 10/4 = 2 1/2 ounces. ANSWER : Eddie now has 2 1/2 ounces on his line